A 610 lines/mm diffraction grating is in an empty aquarium tank. The index of refraction of the glass walls is nglass = 1.50. A helium-neon laser (λ = 633nm) is outside the aquarium. The laser beam passes through the glass wall and illuminates the diffraction grating. What is the first-order diffraction angle of the laser beam? What is the first-order diffraction angle of the laser beam after the aquarium is filled with water ( nwater = 1.33)?
Answer #1

Please rate
if any mistake in this answer please comment i will clarify your doubt . thank you
Answer #2

First-Order Diffraction Angle

Given:

  • Diffraction grating density: 610 lines/mm → Grating spacing d=1610×103m=1.639×106m.

  • Laser wavelength (in air): λair=633nm=633×109m.

  • Refractive indices:

    • Glass (nglass): 1.50

    • Water (nwater): 1.33



Step 1: Empty Aquarium (Glass Walls Only)

  1. Wavelength in glass:

    λglass=λairnglass=633nm1.50=422nm.

  2. Diffraction condition (first order, m=1):

    dsinθ=mλ    sinθ=λglassd.sinθ=422×1091.639×1060.2575.

  3. First-order angle:

    θ=sin1(0.2575)14.9214.9.


Step 2: Aquarium Filled with Water

  1. Wavelength in water:

    λwater=λairnwater=633nm1.33476nm.

  2. Diffraction condition:

    sinθ=λwaterd=476×1091.639×1060.2904.

  3. First-order angle:

    θ=sin1(0.2904)16.8816.9.


answered by: anonymous